Consider the following tridiagonal system of linear equations:

[3 2 0 0 [x1 [12

2 3 2 0 x2 = 17

0 2 3 2 x3 14

0 0 2 3] x4 ] 7]

Solve this system using Thomas algorithm

Let 𝑥(0)T=[0 0 0 0] and solve the system using Jacobi, and Gauss-Seidel for at least five iterations


1
Expert's answer
2021-10-12T02:12:16-0400

Using Thomas algorithm:


"a_1=2\/3,b_1=12\/3=4"

"a_2=2/(3-2\cdot2/3)=6/5,b_2=(17-2\cdot4)/(3-2\cdot2/3)=27/5"

"a_3=2/(3-2\cdot6/5)=10/3,b_3=(14-2\cdot27/5)/(3-2\cdot6/5)=16/3"

"b_4=(7-2\cdot16/3)/(3-2\cdot10/3)=1"


"x_4=b_4=1"

"x_3=b_3-a_3x_4=16\/3-10\/3=2"

"x_2=b_2-a_2x_3=27/5-6\cdot2/5=3"

"x_1=b_1-a_1x_2=4-2\cdot3/3=2"


Using Jacobi method:


"x_1^{(1)}=12\/3=4,x_2^{(1)}=17\/3,x_3^{(1)}=14\/3,x_4^{(1)}=7\/3"

"x_1^{(2)}=1/3(12-2\cdot17/3)=2/9"

"x_2^{(2)}=1/3(17-2\cdot4-2\cdot14/3)=-1/9"

"x_3^{(2)}=1/3(14-2\cdot17/3-2\cdot7/3)=-2/3"

"x_4^{(2)}=1/3(7-2\cdot14/3)=-7/9"

"x_1^{(3)}=1\/3(12+2\/9)=110\/27"

"x_2^{(3)}=1/3(17-2\cdot2/9+2\cdot2/3)=161/27"

"x_3^{(3)}=1/3(14+2/9+2\cdot7/9)=142/27"

"x_4^{(3)}=1/3(7+2\cdot2/3)=25/9"

"x_1^{(4)}=1/3(12-2\cdot161/27)=2/81"

"x_2^{(4)}=1/3(17-2\cdot110/27-2\cdot142/27)=-45/81=-5/9"

"x_3^{(4)}=1/3(14-2\cdot161/27-2\cdot25/9)=-94/81"

"x_4^{(4)}=1/3(7-2\cdot142/27)=-95/81"

"x_1^{(5)}=1/3(12+2\cdot5/9)=118/27=4.37"

"x_2^{(5)}=1/3(17-2\cdot2/81+2\cdot94/81)=1561/243=6.42"

"x_3^{(5)}=1/3(14+2\cdot5/9+2\cdot95/81)=1414/243=5.82"

"x_4^{(5)}=1/3(7+2\cdot94/81)=755/243=3.11"


Using Gauss-Seidel method:

"x_1^{(1)}=12\/3=4"

"x_2^{(1)}=1/3(17-2\cdot4)=3"

"x_3^{(1)}=1/3(14-2\cdot3)=8/3"

"x_4^{(1)}=1/3(7-2\cdot8/3)=5/9"

"x_1^{(2)}=1/3(12-2\cdot3)=2"

"x_2^{(2)}=1/3(17-2\cdot2-2\cdot8/3)=23/9"

"x_3^{(2)}=1/3(14-2\cdot23/9-2\cdot5/9)=90/27=10/3"

"x_4^{(2)}=1/3(7-2\cdot10/3)=1/9"

"x_1^{(3)}=1/3(12-2\cdot23/9)=62/27"

"x_2^{(3)}=1/3(17-2\cdot62/27-2\cdot10/3)=155/81"

"x_3^{(3)}=1/3(14-2\cdot155/27-2\cdot1/9)=90/27=62/81"

"x_4^{(3)}=1/3(7-2\cdot62/81)=443/243"

"x_1^{(4)}=1/3(12-2\cdot155/81)=662/243"

"x_2^{(4)}=1/3(17-2\cdot662/243-2\cdot62/81)=2435/729"

"x_3^{(4)}=1/3(14-2\cdot2435/729-2\cdot443/243)=90/27=1.22"

"x_4^{(4)}=1/3(7-2\cdot1.22)=1.52"

"x_1^{(5)}=1/3(12-2\cdot2435/729)=1.77"

"x_2^{(5)}=1/3(17-2\cdot1.77-2\cdot1.22)=3.67"

"x_3^{(5)}=1/3(14-2\cdot3.67-2\cdot1.52)=90/27=1.21"

"x_4^{(5)}=1/3(7-2\cdot1.21)=1.53"